# RankCLIP: Ranking-Consistent Language-Image Pretraining

### v2

@@ 4.4 Linear Probing
We also evaluate whether the introduced ranking consistency retains its advantages when supplemented with additional in-domain supervision. Specifically, we use linear probing, where the pretrained encoders from CLIP, ALIP, and RankCLIP remain unchanged while a logistic regression classifier is trained on domain-specific datasets. We evaluate on a suite of 11 standard image classification datasets as our in-domain datasets, which include CIFAR-10, CIFAR-100, Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Fine-Grained Visual Classification of Aircraft (FGVG-Aircraft) (Maji et al., 2013), Food101 (Bossard et al., 2014), German Traffic Sign Detection Benchmark (GTSDB) (Stallkamp et al., 2012), ImageNet1K (Deng et al., 2009; Russakovsky et al., 2015), OxfordPets (Parkhi et al., 2012), Stanford Sentiment Treebank v2 (SST2) (Socher et al., 2013), STL-10 (Coates et al., 2011), and Street View House Numbers (SVHN) (Netzer et al., 2011) dataset.
Table 3 indicates that RankCLIP consistently outperforms CLIP, with relative improvements ranging from +0.2% to +11.5% and an average accuracy increase of +2.31%. When compared to ALIP, RankCLIP also shows better performance on average, though the gains are relatively modest.
4.5 Zero-shot Image-text Retrieval
In the final part of our experiments, we assess RankCLIP on zero-shot cross-modal retrieval tasks (image-to-text and text-to-image) using the Flickr30k (Plummer et al., 2015) and MSCOCO (Lin et al., 2014) datasets. As shown in Table 4, RankCLIP generally outperforms the two baseline methods, though improvements are less significant compared to earlier results in Table 1, 2 and 3. The relatively modest gains in retrieval tasks may stem from the complex requirements of discerning image-text similarities across varying resolutions and object details, a significant departure from the simpler demands of image classification tasks. Despite this, the overall improvement highlights RankCLIP’s advantage, thanks to the deeper insights provided by ranking consistency in the language-image training process.
@@ 5 Ablation Studies
| CIFAR-10 | CIFAR-100 | ImageNet1K | |||||||
| Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | Top1 | Top3 | Top5 | |
| CLIP | 36.35% | 70.28% | 85.02% | 12.22% | 24.93% | 33.56% | 12.08% | 21.86% | 27.48% |
| RankCLIP | 37.03% | 67.67% | 83.09% | 13.98% | 27.70% | 36.17% | 17.02% | 28.44% | 33.99% |
| 37.47% | 69.89% | 84.53% | 13.89% | 27.34% | 35.90% | 16.66% | 27.63% | 33.15% | |
| 28.26% | 59.65% | 75.45% | 13.29% | 26.85% | 34.71% | 16.98% | 28.25% | 33.90% | |
5.1 Ablation on Loss Components
To further assess the effectiveness of the proposed ranking consistency, we developed two variants of RankCLIP: , focusing solely on cross-modal consistency with , and , emphasizing in-modal consistency with . Both models underwent the same pretraining as outlined in Appendix B and were tested in a zero-shot classification experiment on ImageNet1K as in §4.2. The results are shown in Table 5, with bold font indicating the best performance, and blue color representing the second best results. We can see that, while RankCLIP achieves the best performance, both and demonstrate notable improvements over CLIP. Interestingly, matches the performance of , highlighting the often-underestimated value of in-modal consistency in enhancing model effectiveness.
5.2 Ablation on Data Sizes
To evaluate the scalability of RankCLIP, we trained both CLIP and RankCLIP using 500K, 750K, 1M, and 3M text-image pairs from the CC3M dataset following the same procedures in §B. Fig. 3 left presents the zero-shot top-1 classification accuracy on ImageNet1K, where RankCLIP consistently outperforms CLIP. Notably, it shows a greater performance increase as dataset size grows from 1M to 3M pairs, suggesting RankCLIP’s superior scalability, a critical attribute for language-image pretraining. Furthermore, Fig. 3 right illustrates RankCLIP’s robustness across different dataset sizes. The horizontal axis shows the top-1 accuracy on standard ImageNet1K, and the vertical axis on ImageNet1K-R, with a black diagonal line () representing ideal robustness. Any deviation below this line indicates reduced robustness. RankCLIP consistently stays well above both the red baseline, which reflects typical in-distribution to out-of-distribution generalization (Miller et al., 2021), and close to the ideal line, demonstrating exceptional robustness to distribution shifts.
@@ 6.1 Modality Gap
In this section, we analyze the modality gaps of CLIP and our proposed RankCLIP by visualizing 250 text-image pair embeddings, reduced to two dimensions using UMAP (McInnes et al., 2018), and complement this with a histogram of the gaps. Modality gap (Liang et al., 2022) refers to a geometric phenomenon observed in the representation spaces of multimodal models, where different data modalities (like images and texts) are embedded at a noticeable distance from each other, rather than being uniformly distributed as ideally expected. This gap, inherent from initialization and preserved during the contrastive learning process like in CLIP, poses a challenge in language-image pretraining by impacting joint data modeling and understanding. Recent studies (Srivastava and Sharma, 2024; Kumar and Marttinen, 2024; Oh et al., 2024) suggest that reducing this gap could enhance multimodal representations and downstream task performance. The results shown in Fig. 4 indicate that RankCLIP exhibits a significantly smaller modality gap than CLIP, demonstrating that our ranking consistency approach effectively enhances understanding of text-image semantics.
@@ 6.2 Alignment and Uniformity
Besides alleviating modality gap, it is also commonly believed that a successful contrastive learning method should as well ensure a broad and uniform distribution covering an hypersphere in space (Wang and Isola, 2020). These two goals, characterized as similarity and uniformity, can be assessed with alignment and uniformity scores, respectively. More specifically, following Goel et al. (2022) and notations defined in §3, we calculate the alignment score , and uniformity score to be:
where is the total number of text-image pairs. Essentially, represents the averaged cosine similarity between text and image embeddings, and averages the dissimilarity measures (exponentiated negative dot products) between all unique pairs of text-image embeddings in the dataset, quantifying how evenly these embeddings are distributed.
A high alignment score represents a strong correlation or similarity between pairs of text-image embeddings, indicating that the images and their textual descriptions are closely aligned in the embedding space. Conversely, a high uniformity score suggests that embeddings are not uniformly distributed; they may be clustering together or not utilizing the embedding space efficiently, which can indicate redundancy in the representations or a lack of diversity. A low uniformity score, on the other hand, suggests that the embeddings are well spread out across the space, indicating a diverse and efficient use of the embedding space, which is generally desirable for tasks like retrieval, where a wide coverage of possible queries are preferred.
As shown in Table 6, we observe that, although CLIP learns representations that are better aligned, as evidenced by its top-ranking alignment scores, these representations fail to achieve uniform distribution across the hypersphere, as highlighted by its significantly higher absolute uniformity scores. On the other hand, RankCLIP, along with two of its ablated version, and , presents much better balance between alignment and uniformity, which results in improved downstream task performance as illustrated in previous experiments as well as in the representative ZS-Top1 results in Table 6. We also find the results to be informative on a higher level where it indicates that optimizing contrastive learning towards single objective such as alignment or uniformity would not intuitively result in higher downstream task performance.
| CIFAR-10 | CIFAR-100 | ImageNet1K | |||||||
| ZS-Top1 | ZS-Top1 | ZS-Top1 | |||||||
| CLIP | 0.40 | -0.35 | 36.35% | 0.42 | -0.35 | 12.22% | 0.44 | -0.29 | 12.08% |
| RankCLIP | 0.23 | -0.17 | 37.03% | 0.26 | -0.16 | 13.98% | 0.33 | -0.11 | 17.02% |
| 0.24 | -0.16 | 37.47% | 0.26 | -0.15 | 13.89% | 0.32 | -0.10 | 16.66% | |
| 0.18 | -0.12 | 28.26% | 0.18 | -0.10 | 13.29% | 0.26 | -0.09 | 16.98% | |
@@ 7 Conclusion
Limitations
This work has two main limitations. First, due to computational constraints, we were not able to perform more baseline comparisons by pretraining additional existing methods. And second, we were unable to scale the pretraining process to include larger datasets, such as those used in OpenCLIP (Ilharco et al., 2021).
@@
Appendix A CLIP Preliminaries
CLIP (Radford et al., 2021) has been a prominent method for learning detailed multimodal representations through the alignment of images and texts. Given a set of image-text pairs, where denotes an image and is the corresponding text, the goal is to learn representations that map semantically similar images and texts closer in the embedding space, while dissimilar pairs are distanced apart. More specifically, the foundational CLIP model employs two encoders: an image encoder that processes raw images into visual embeddings and a text encoder which encodes textual data into text embeddings. Then both the text and visual features are projected to a latent space with identical dimension. Formally, the embeddings for a text-image pair are denoted as and , respectively. The embeddings are then normalized to lie on an unit hypersphere by enforcing -norm constraint:
@@ Appendix A CLIP Preliminaries
To align the image and text representations, a contrastive loss function, typically a variant of the InfoNCE loss (Oord et al., 2018), which optimizes the similarity of the matched pair against unmatched pairs, is utilized, i.e.:
@@
Appendix B Training Procedures
B.1 Implementation Details
For CLIP (Radford et al., 2021), we use the official implementation released by OpenAI111CLIP repository on GitHub: https://github.com/openai/CLIP.. And for ALIP (Yang et al., 2023), we also use the official implementation released by the paper authors222ALIP repository on GitHub: https://github.com/deepglint/ALIP.. As the proposed RankCLIP essentially shares the same model architecture (separate vision, text encoders, projection layer, and a classification head) as CLIP, we build upon the CLIP code repository for our model construction333RankCLIP repository will be released upon acceptance.. We set the scaling parameters for cross-modal () and in-modal () ranking consistency to 1/16 and 1/16 respectively throughout all the experiments unless otherwise noted. All CLIP, ALIP and RankCLIP models are initialized from scratch without loading any existing weights. And the embedding sizes for both modalities all project to 1024 across the three models.
B.2 Training Parameters
Following CLIP (Radford et al., 2021), we adopt the ResNet-50 (He et al., 2016) and transformer architectures (Devlin et al., 2018) for image and text encoding, respectively. Training is conducted from scratch over 64 epochs using a single NVIDIA A100 GPU, with a batch size of 512, an initial learning rate of 0.0005 employing cosine scheduling, and 10,000 warm-up steps.

### v3

@@ 4.5 Linear Probing
We further assess whether the advantages of ranking consistency persist when supplemented with in-domain supervision. Specifically, we apply linear probing, where pretrained encoders remain fixed while a logistic regression classifier is trained on domain-specific datasets. We evaluate on 10 standard image classification benchmarks, including CIFAR-10, CIFAR-100, DTD [9], FGVG-Aircraft [33], Food101 [1], GTSDB [52], OxfordPets [40], SST2 [50], STL-10 [10], and SVHN [37]. As shown in Table 3, RankCLIP achieves the highest average accuracy, demonstrating that ranking consistency enhances generalization even with additional in-domain supervision.
@@ 5 Ablation Studies
5.1 Different Weights of RankCLIP Loss
In Eq. (8), we define the RankCLIP loss as a linear combination of the original contrastive loss and the in-modal and cross-modal ranking consistency losses, weighted by and , respectively. In §3.5, we introduce a training strategy that adaptively adjusts and at different stages of training. In this section, we analyze the impact of these weights and demonstrates that adaptive weighting further enhances RankCLIP ’s performance. Notice that all model variants follow the same pretraining setup detailed in Appendix A. As shown in Fig. 3, RankCLIP outperforms CLIP even with fixed and , highlighting the effectiveness of ranking consistency. And the adaptive weighting strategy further boosts accuracy by preventing the under-utilization of ranking consistency at low weights and avoiding disruptions to contrastive learning at high weights.
Method | Vision Backbone | ImageNet1K | MSCOCO | Linear Probing Avg. Acc. | |||||||
Image-to-Text Retrieval | Text-to-Image Retrieval | ||||||||||
Top-1 | Top-3 | Top-5 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | |||
CLIP [43] | RN50 | 21.6% | 36.9% | 44.9% | 15.6% | 36.4% | 48.4% | 6.7% | 15.2% | 20.1% | 64.2% |
RankCLIP | 30.9% | 49.4% | 57.6% | 19.5% | 42.6% | 54.8% | 7.5% | 16.2% | 21.6% | 68.9% | |
CLIP [43] | ViT-B/32 | 20.7% | 35.0% | 42.4% | 11.9% | 29.4% | 40.8% | 5.1% | 12.9% | 17.9% | 60.7% |
RankCLIP | 26.2% | 41.4% | 48.9% | 13.8% | 33.8% | 45.9% | 6.0% | 13.6% | 18.6% | 61.3% |
5.2 Different Data Sizes
To assess the scalability of RankCLIP, we trained both CLIP and RankCLIP on 500k, 750k, 1M, and 3M text-image pairs from CC3M, as well as 15M pairs from YFCC15M, following the procedure in Appendix A. Fig. 4 compares their performance on zero-shot top-1 classification accuracy for ImageNet1K (left) and averaged linear probing results (middle), where RankCLIP consistently outperforms CLIP. Full, non-averaged linear probing results are provided in Appendix 7. Notably, the performance gains of RankCLIP become more pronounced as dataset size scales from 1m to 15m, highlighting its superior scalability, which is critical for language-image pretraining.
Fig. 4 (right) further illustrates RankCLIP’s robustness across different dataset sizes. The horizontal axis represents top-1 accuracy on standard ImageNet1K, while the vertical axis shows accuracy on ImageNet1K-R. The black diagonal () denotes ideal robustness, with deviations below it indicating degradation under distribution shifts. RankCLIP remains well above the red baseline, which represents typical in-distribution to out-of-distribution generalization [35], and stays close to the ideal line, demonstrating strong robustness.
5.3 Different Backbones: RN50 vs. ViT
We compare RankCLIP and CLIP across ResNet-50 (RN50) and ViT-B/32 backbones to assess its generalization. As shown in Table 4, RankCLIP consistently outperforms CLIP across all tasks. Specifically, for zero-shot classification on ImageNet1K, RankCLIP improves top-1 accuracy by +9.3% with RN50 and +5.5% with ViT-B/32, demonstrating stronger feature learning. In cross-modal retrieval, RankCLIP achieves +3.9% in image-to-text and +0.8% in text-to-image with RN50, with smaller but consistent gains for ViT-B/32. RankCLIP also improves linear probing accuracy by +4.7% with RN50 and +0.6% with ViT-B/32, confirming its advantage in representation learning. While both architectures benefit, RN50 sees the largest gains, suggesting that ranking consistency particularly enhances hierarchical feature extraction in CNNs.
@@ 6.1 Modality Gap
We analyze the modality gaps of CLIP and RankCLIP by visualizing 250 text-image pair embeddings, reduced to two dimensions using UMAP [34], and presenting a histogram of the gaps. The modality gap [29] refers to the separation between text and image embeddings in multimodal models, hindering joint representation learning. This gap, inherent from initialization and reinforced by contrastive learning in CLIP, challenges effective language-image modeling. Recent studies [51, 27, 38] suggest that reducing this gap improves multimodal representations and downstream performance. As shown in Fig. 6, RankCLIP exhibits a significantly smaller modality gap than CLIP, demonstrating that our ranking consistency approach effectively enhances text-image alignment.
@@ 6.2 Alignment and Uniformity
Beyond reducing modality gap, effective contrastive learning should ensure a broad and uniform distribution over a hypersphere [56]. These objectives – similarity and uniformity – are quantified by alignment and uniformity scores, respectively. Following Goel et al. [19] and the notations in §3, we compute the alignment score and uniformity score as:
where is the number of text-image pairs.
captures the average cosine similarity between corresponding text and image embeddings, while measures how evenly embeddings are spread across the space. High alignment indicates strong correlation between paired embeddings, whereas low uniformity suggests diverse and efficient embedding distribution—desirable for tasks like retrieval. As shown in Table 5, CLIP achieves stronger alignment but suffers from poor uniformity, leading to redundant representations. On the other hand, RankCLIP along with two of its ablated version, and presents much better balance between alignment and uniformity. These results further suggest that optimizing solely for alignment or uniformity does not necessarily translate to better task performance.
CIFAR-10 | CIFAR-100 | ImageNet1K | ||||
CLIP | 0.28 | -0.19 | 0.28 | -0.18 | 0.33 | -0.19 |
RankCLIP | 0.23 | -0.14 | 0.26 | -0.13 | 0.29 | -0.13 |
0.23 | -0.13 | 0.24 | -0.12 | 0.29 | -0.14 | |
0.25 | -0.15 | 0.28 | -0.14 | 0.32 | -0.17 |
6.3 Qualitative Examples
Class activation maps.
To further examine the effects of ranking consistency, we visualize class activation maps (CAMs) [47] for RankCLIP and CLIP in Fig. 7. The results show that RankCLIP consistently attends to more semantically relevant regions in the images. For example, when given the caption ‘An airport filled with planes sitting on tarmacs’, CLIP mistakenly highlights surrounding areas, whereas RankCLIP focuses precisely on the planes. Similar improvements are observed across other examples, demonstrating that RankCLIP better aligns textual descriptions with visual concepts. This suggests that ranking consistency enhances fine-grained feature learning, leading to more localized and accurate visual grounding.
Text-to-image retrieval.
Fig. 5 compares text-to-image retrieval results from RankCLIP and CLIP. Given a query, we display the top-ranked images retrieved by each model. RankCLIP consistently retrieves more semantically aligned images, even beyond the correct reference image. For instance, in the example of ’A cute cat laying down in a sink’, correctly identifies a cat in a sink, whereas CLIP misidentifies it due to the visual similarity between sinks and toilets. This demonstrates RankCLIP ’s ability to capture fine-grained semantic distinctions, reinforcing its advantage in retrieval tasks that demand precise understanding.
@@ 7 Conclusion
Appendix A Training Procedures
A.1 Implementation Details
For CLIP [43], we use the official implementation released by OpenAI111CLIP repository on GitHub: https://github.com/openai/CLIP.. And for ALIP [58], we also use the official implementation released by the paper authors222ALIP repository on GitHub: https://github.com/deepglint/ALIP.. As the proposed RankCLIP essentially shares the same model architecture (separate vision, text encoders, projection layer, and a classification head) as CLIP, we build upon the CLIP code repository for our model construction333RankCLIP repository will be released upon acceptance.. We set the scaling parameters for cross-modal () and in-modal () ranking consistency to 1/16 and 1/16 respectively throughout all the experiments unless otherwise noted. All CLIP, ALIP and RankCLIP models are initialized from scratch without loading any existing weights. And the embedding sizes for both modalities all project to 1024 across the three models.
A.2 Training Parameters
Following CLIP [43], we adopt the ResNet-50 [21] and transformer architectures [13] for image and text encoding, respectively. Training is conducted from scratch over 64 epochs using a single NVIDIA A100 GPU, with a batch size of 512, an initial learning rate of 0.0005 employing cosine scheduling, and 10,000 warm-up steps.
A.3 Training Time Consumption
we conducted the experiments using the same hardware specifications. The table below shows the time consumption for training our RankCLIP and CLIP models with 50K samples from CC3M using a single NVIDIA A100 GPU.
Time consumption | Dataset size | epochs | batch_size | model_name | |
CLIP | 1d 2h 54m 48s | 50K | 64 | 512 | RN50 |
RankCLIP | 1d 1h 4m 23s | 50K | 64 | 512 | RN50 |
As shown in the table, the difference in time consumption is negligible. Interestingly, our method is slightly faster than CLIP, but we think it may be attributed to hardware optimizations or variance.
Appendix B CLIP Preliminaries
CLIP [43] has been a prominent method for learning detailed multimodal representations through the alignment of images and texts. Given a set of image-text pairs, where denotes an image and is the corresponding text, the goal is to learn representations that map semantically similar images and texts closer in the embedding space, while dissimilar pairs are distanced apart. More specifically, the foundational CLIP model employs two encoders: an image encoder that processes raw images into visual embeddings and a text encoder which encodes textual data into text embeddings. Then both the text and visual features are projected to a latent space with identical dimension. Formally, the embeddings for a text-image pair are denoted as and , respectively. The embeddings are then normalized to lie on an unit hypersphere by enforcing -norm constraint:
@@ Appendix B CLIP Preliminaries
To align the image and text representations, a contrastive loss function, typically a variant of the InfoNCE loss [39], which optimizes the similarity of the matched pair against unmatched pairs, is utilized, i.e.:
@@
B.1 Additional Experiments
We conduct the linear probing experiment under different training datasize from 3m to 15m as shown in 7.
Data Size | Method | Model Type |
CIFAR-10 |
CIFAR-100 |
DTD |
FGVGAircraft |
Food101 |
GTSRB |
OxfordPets |
SST2 |
STL10 |
SVHN |
CLIP | RN50 | 80.12% | 58.50% | 57.18% | 39.75% | 59.14% | 72.41% | 61.73% | 54.48% | 86.01% | 58.92% | |
3m | RankCLIP | RN50 | 78.29% | 56.24% | 57.82% | 39.30% | 58.63% | 74.13% | 64.35% | 55.02% | 86.69% | 60.68% |
CLIP | ViT-B/32 | 77.60% | 56.15% | 43.19% | 22.59% | 39.72% | 62.05% | 40.39% | 50.96% | 78.99% | 50.53% | |
RankCLIP | ViT-B/32 | 78.42% | 56.64% | 42.39% | 23.43% | 40.19% | 60.63% | 40.56% | 53.32% | 79.60% | 47.72% | |
CLIP | RN50 | 78.81% | 56.32% | 61.49% | 25.83% | 61.64% | 68.76% | 60.37% | 55.57% | 89.82% | 47.99% | |
15m | RankCLIP | RN50 | 83.27% | 62.96% | 65.96% | 32.19% | 68.11% | 74.25% | 67.40% | 56.34% | 94.20% | 53.06% |
CLIP | ViT-B/32 | 82.97% | 62.55% | 49.47% | 24.48% | 52.46% | 63.55% | 50.78% | 52.66% | 87.14% | 46.38% | |
RankCLIP | ViT-B/32 | 82.79% | 59.89% | 52.50% | 23.94% | 56.44% | 61.58% | 52.98% | 53.60% | 89.01% | 42.16% |